# Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models

### v2

@@ 3.4 FILA: a Novel LoRA Initialization for LLM Unlearning
Motivation. Biderman et al. (2024) discovered that, while it typically does not surpass full fine-tuning in performance, LoRA also induces less forgetting than full fine-tuning in domain adaptation scenarios. While this stability is beneficial to retaining knowledge of ${\mathcal{D}}_{r}$ during unlearning, it also imposes a strong burden under the strict objective to remove knowledge of ${\mathcal{D}}_{f}$ completely. As a result, vanilla LoRA unlearning requires large number of model updates through ${\mathcal{D}}_{f}$ for successful unlearning, which leads to significant deterioration in downstream performance (as shown in Figure 2). Drawing inspiration from PiSSA (Meng et al., 2024), we conjecture that initializing LoRA adapters to contain parameters that are relatively more important to ${\mathcal{D}}_{f}$ than to ${\mathcal{D}}_{r}$ would reinforce the model’s plasticity to forgetting ${\mathcal{D}}_{f}$, thereby accelerating the rate of unlearning and minimizing catastrophic forgetting on ${\mathcal{D}}_{r}$. In order to quantify the importance of parameters to particular data points, we leverage Fisher information matrices.
Parameter Importances via Fisher Information. The Fisher information matrix ${\bm{F}}_{\theta}$ captures the amount of information a dataset ${\mathcal{D}}$ provides on model parameters $\theta$. More concretely, ${\bm{F}}_{\theta}$ is computed as the second cross-moments of first partial derivatives of the log-likelihood of ${\mathcal{D}}$ (left of Eq. 2). However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}$ instead. In the context of LLMs, this can be computed as:
@@
Let $\hat{{\bm{F}}}_{\bm{W}}^{f}\coloneqq\hat{{\bm{F}}}_{\bm{W}}({\mathcal{D}}_{f})$ denote the empirical Fisher information matrix of the target parameter ${\bm{W}}$ measured using the forget set ${\mathcal{D}}_{f}$ (resp. $\hat{{\bm{F}}}_{\bm{W}}^{r}$ using the retain set ${\mathcal{D}}_{r}$). Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times r}$ as an importance metric to identify parameters that are important exclusively for ${\mathcal{D}}_{f}$ and not for ${\mathcal{D}}_{r}$. While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus use $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$ rather than $\hat{{\bm{F}}}_{\bm{W}}^{f}$.
@@
with $\mathds{1}\in\mathbb{R}^{k}$ and $\texttt{diag}(\cdot)$ indicating the all-one vector and the vector diagonalization function, respectively. Unlike general WLRA, this row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\text{diag}(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{-1}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
Given this solution, we use ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights. To ensure that the model behavior remains the same after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}^{*}{\bm{A}}^{*}$. Intuitively, our Fisher-weighted Initialization of Low-rank Adapters (FILA) extracts parameters that are important for generating ${\mathcal{D}}_{f}$, but not for generating ${\mathcal{D}}_{r}$, such that LoRA tuning can be focused on erasing knowledge relevant to ${\mathcal{D}}_{f}$ while keeping information regarding ${\mathcal{D}}_{r}$.
@@ 3.5 Final Loss Function for LLM Unlearning
In summary, we perform unlearning on the model $\bm{\Theta}=\bm{\theta}\cup\bm{\theta}_{\text{FILA}}$, consisted of original pretrained weights $\bm{\theta}$ and the FILA-initialized low-rank adapter weights for each linear layer $\bm{\theta}_{\text{FILA}}=\{{\bm{A}}^{*}_{\ell},{\bm{B}}^{*}_{\ell}\}_{\ell=1}^{L}$, where $L$ represents the number of layers tuned via LoRA. Additionally, we incorporate GD, which utilizes the auxiliary retain set $\mathcal{D}_{r}$. The final loss function using both the proposed IHL and FILA is defined as follows:
@@
In practice, training (unlearning) for the LLM model is conducted by minimizing Eq.3 through stochastic gradient descent.
@@ 4 Experiments
Table 1: Evaluation results on LM-Evaluation, Dialogue datasets before and after unlearning samples from the TDEC dataset. For all runs under LoRA, we use rank 16 and GD is used as the baseline. The "+IHL" results refer to experiments where NCE loss within GD is replaced with our proposed IHL (i.e., Equation 3), and "+FILA" indicates results where FILA is used to initialize LoRA.
@@
Before - - 30.9 77.4 43.4 9.4 17.8
GA 100.0 17.2 1.0 27.4 39.9 2.6 577.8
GPT-Neo GD 4.6 0.7 24.9 42.4 5.9 54.2
125M LoRA 1.6 8.6 0.3 20.6 40.8 2.5 129.4
+ IHL 11.4 0.4 22.7 41.9 6.0 32.9
     + FILA 6.0 0.3 23.9 42.2 10.1 24.0
Before - - 67.6 92.2 49.8 11.5 11.5
GA 100.0 13.8 1.9 30.4 49.7 8.5 15.8
GPT-Neo GD 12.8 2.2 30.9 48.4 12.7 10.8
1.3B LoRA 0.8 19.3 1.7 31.4 45.0 9.7 31.8
+ IHL 20.0 1.7 44.6 47.1 10.2 14.9
     + FILA 13.0 0.5 29.6 48.3 12.1 14.7
Before - - 70.4 93.4 52.3 11.5 10.4
GA 100.0 10.8 1.6 31.0 51.9 11.1 17.9
GPT-Neo GD 8.0 0.7 28.3 44.0 12.7 17.9
2.7B LoRA 0.7 14.0 0.1 20.4 45.9 6.7 61.1
+ IHL 17.8 0.0 26.7 49.6 8.5 22.2
     + FILA 10.3 0.1 28.5 49.6 10.7 16.0
@@ 4.1 Training Data Extraction Challenge
Experimental Setup.    The Training Data Extraction Challenge (TDEC) dataset  [^1] The dataset was originally published as part of a competition held at SaTML 2023: https://github.com/google-research/lm-extraction-benchmark consists of 20k examples from the Pile dataset (Gao et al., 2020) found to be easily extractable from a pretrained LLM. For each experiment, we randomly sample 32 sequences with 200 tokens to consist the forget set ${\mathcal{D}}_{f}$. For the retain set ${\mathcal{D}}_{r}$, we use the subset of WikiText (Merity et al., 2017) as it contains factual world knowledge that we wish to maintain after unlearning. We consider GPT-Neo 125M, 1.3B, and 2.7B pretrained on the Pile dataset as our base models, and unlearn ${\mathcal{D}}_{f}$ using five different forget sets. For this experiment, we use a fixed learning rate of 2e-4 and use LoRA adapters with rank $r=\{4,8,16,32\}$. For reasons we illustrate later in §4.2, we choose to apply LoRA on query and value layers in the attention module and two linear layers within feed-forward layers.
@@
Figure 3: Results from unlearning examples in the TDEC dataset from the GPT-Neo LLMs with three different sizes. Each row represents the performance averaged across datasets within each set of LLM capability tests: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better). The circles and crosses represent successful and unsuccessful attempts, respectively, of unlearning a particular forget set ${\mathcal{D}}_{f}$. Solid lines indicate the performance of different methods averaged only across successful unlearning trials. The dashed lines indicate the base model performance prior to unlearning. Unlearning with GD leads to significant loss in performance and also frequently fails to unlearn effectively even with large LoRA ranks. Replacing the negative cross-entropy (NCE) loss with our hinge loss (IHL) boosts retention of reasoning and generation capabilities, but still fails to unlearn in multiple cases. Using NCE loss with FILA initialization notably increases the rate of unlearning success, but at the cost of losing overall performance. Using both IHL and FILA best minimizes post-unlearning performance degradation in all three aspects.
@@
Results.   Table 1 shows evaluation results using a fixed LoRA rank of 16 and Figure 3 shows analogous results using a different LoRA ranks. Our key findings are as follows. First, using GD not only meets the forgetting criteria in fewer epochs across all model sizes but it also preserves previously acquired knowledge (i.e., performance on Reasoning, Dialogue, and Pile) better than GA. Mainly for 125M and 1.3B models, GA causes a significant decline in generative performance, whereas GD partially mitigates this decline and improves both Dialogue F1 scores and the Pile perplexity. Second, we find that simply applying GD with LoRA fails to achieve successful unlearning across all model sizes. Second, we find that imply running GD with LoRA significantly increases the number of epochs required for successful unlearning. While enjoying great parameter-efficiency by tuning only about 0.7% to 1.6% of the total parameters when using LoRA, its application to unlearning with GD results in significant loss in overall language capability, especially on generative tasks. Lastly, using our proposed IHL in conjunction with FILA achieves the forgetting criteria in fewest epochs while best preserving LLM performance. Although IHL with LoRA alone retains knowledge better than the baseline, it requires more epochs to meet the forgetting criteria. This is resolved when used together with FILA, and the number of required epochs is significantly reduced, reducing catastrophic forggetting of previously acquired knowledge even further.
Figure 4: Results from unlearning examples from TDEC dataset using LoRA with rank 32 to adapt sets of layers on GPT-Neo 2.7B. The marker shapes and colors are used similarly as in Figure 3. Based on the rate of unlearning success, tuning FFN layers (e.g., FFN, QVFFN) is more receptive to targeted knowledge removal compared to tuning attention layers (e.g., QV, QKVO).
@@ 4.3 Task of Fictitious Unlearning
Experimental Setup.  The Task of Fictitious Unlearning (TOFU) benchmark (Maini et al., 2024) is a synthetic dataset containing 20 question-answer pairs for each of 200 fictitious author profiles generated by GPT-4. The TOFU evaluation pipeline first finetunes a pretrained LLM on all QA pairs. Given this finetuned LLM that serves as our base model, our task is to unlearn all information regarding 1%, 5%, or 10% of the authors from the model. Note that we can obtain reference models finetuned only on the retain set (QA-pairs on 99%, 95%, or 90% of authors), with which we evaluate the Forget Quality of unlearned models by measuring the p-value from a Kolmogorov-Smirnov test. A high p-value indicates high distributional similarity between the unlearned model and the reference model, thus implying strong forgetting. To evaluate how well the model retains other information outside the forget set, we measure the Model Utility as the aggregated model performance. Further details on the dataset and evaluation pipeline can be found in  Maini et al. (2024).
Following the original paper of TOFU, we prepare two base models by finetuning Phi-1.5B and Llama2-7B on TOFU for 5 epochs with learning rates 2e-5 and 1e-5, respectively. We then unlearn using two baselines (GA and GD) and our two methods (IHL and IHL+FILA) using LoRA adapters of rank 4, 8, 16, or 32. For unlearning, we use a learning rate of 2e-4 if our base model is from Phi-1.5B and 1e-4 for Llama2-7B. All training procedures run 5 epochs with an effective batch size of 32 using the AdamW optimizer (Loshchilov & Hutter, 2019).
@@
[…]
@@
With regards to our proposed method, we find that replacing the NCE loss in GD with our IHL better retains model utility across all LoRA ranks and forget set sizes, as curves are more aligned straight-up towards the reference point with negligible shift in model utility. This stability comes at the cost of unlearning efficiency, however, as the rate at which the LLM forgets ${\mathcal{D}}_{f}$ is slower with IHL due to IHL decreasing the likelihood of the unwanted token by increasing the likelihood of only one other most-possible token each time in a controlled manner. Nonetheless, initializing LoRA adapters with FILA largely alleviates this issue and enhances unlearning efficiency of IHL by focusing gradient updates on parameters important to generating ${\mathcal{D}}_{f}$.

### v3

@@ 3.4 FILA: a Novel LoRA Initialization for LLM Unlearning
Motivation. While IHL effectively stabilizes the unlearning process by reducing the likelihood of unwanted tokens in a controlled manner, we empirically find that combining IHL with LoRA naïvely leads to requiring large number of unlearning iterations to fully forget samples in ${\mathcal{D}}_{f}$. Naturally, an increasing number of updates with ${\mathcal{D}}_{f}$ comes with the risk of overfitting, which for unlearning, may result in significant loss of knowledge on ${\mathcal{D}}_{r}$. We hypothesize that this is due to the random initialization of LoRA weights together with its low-rank structure imposing too strong a regularization to handle precise gradients needed for proper minimization of IHL (Biderman et al., 2024). Drawing inspiration from PiSSA (Meng et al., 2024), we therefore aim to accelerate optimization of IHL under LoRA by extracting weights relatively more important to ${\mathcal{D}}_{f}$ than to ${\mathcal{D}}_{r}$ from each pretrained weight, then using them to initialize LoRA weights prior to unlearning. We conjecture that this approach reinforces the model’s plasticity to forget ${\mathcal{D}}_{f}$ as well as its stability on keeping knowledge on ${\mathcal{D}}_{r}$. The remainder of this section presents how parameter importances are measured via Fisher information, and how low-rank adapter weights are initialized based on the measured importances.
Parameter Importances via Fisher Information. The Fisher information matrix ${\bm{F}}_{\theta}({\mathcal{D}})$ is defined as the variance of the partial derivative of the log-likelihood of data ${\mathcal{D}}$ with respect to the model parmeter $\theta$ (left of Equation 2). Intuitively, the matrix can be considered a measurement on how much the model output changes following a small change on its parameter weight. However, as marginalizing across the space of ${\mathcal{D}}$ is intractable, many works in continual learning (Kirkpatrick et al., 2017) and model compression (Hsu et al., 2022) literature have thus used the empirical Fisher information $\hat{{\bm{F}}}_{\theta}({\mathcal{D}})$ instead. In the context of LLMs, this can be computed as:
@@
Let $\hat{{\bm{F}}}_{\bm{W}}^{f}\coloneqq\hat{{\bm{F}}}_{\bm{W}}({\mathcal{D}}_{f})$ denote the empirical Fisher information matrix of the target parameter ${\bm{W}}$ measured using the forget set ${\mathcal{D}}_{f}$ (resp. $\hat{{\bm{F}}}_{\bm{W}}^{r}$ using the retain set ${\mathcal{D}}_{r}$). Then, we use the relative Fisher information $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\coloneqq\hat{{\bm{F}}}_{\bm{W}}^{f}/\hat{{\bm{F}}}_{\bm{W}}^{r}\in\mathbb{R}^{d\times k}$ as an importance metric to identify parameters that are important exclusively for ${\mathcal{D}}_{f}$ and not for ${\mathcal{D}}_{r}$. While generating ${\mathcal{D}}_{f}$ involves extracting memorized information on ${\mathcal{D}}_{f}$ as well as composing linguistically fluent outputs, we only wish to adjust parameters responsible for the former and thus use $\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}$ rather than $\hat{{\bm{F}}}_{\bm{W}}^{f}$.
@@
with $\mathds{1}\in\mathbb{R}^{k}$ and $\texttt{diag}(\cdot)$ indicating the all-one vector and the vector diagonalization function, respectively. Unlike general WLRA, this row-wise WLRA problem has a closed-form solution, which can be obtained by applying rank-$r$ SVD to decompose $\texttt{diag}((\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{\frac{1}{2}}){\bm{W}}={\bm{U}}{\bm{S}}{\bm{V}}^{T}$ and computing ${\bm{B}}^{*}=(\hat{{\bm{F}}}_{\bm{W}}^{\text{rel}}\mathds{1})^{-\frac{1}{2}}{\bm{U}}{\bm{S}}^{\frac{1}{2}}$ and ${\bm{A}}^{*}={\bm{S}}^{\frac{1}{2}}{\bm{V}}^{T}$.
Given this solution, we use ${\bm{B}}^{*}$ and ${\bm{A}}^{*}$ as initial LoRA weights. To ensure that the model behavior remains the same after LoRA initialization, the base layers are also updated with ${\bm{W}}^{*}={\bm{W}}-{\bm{B}}^{*}{\bm{A}}^{*}$. Intuitively, our Fisher-weighted Initialization of Low-rank Adapters (FILA) extracts parameters that are important for generating ${\mathcal{D}}_{f}$ but not for generating ${\mathcal{D}}_{r}$, such that LoRA tuning can be focused on erasing knowledge relevant to ${\mathcal{D}}_{f}$ while keeping information regarding ${\mathcal{D}}_{r}$.
@@ 3.5 Final Loss Function for LLM Unlearning
In summary, we perform unlearning on the model $\bm{\Theta}=\bm{\theta}\cup\bm{\theta}_{\text{FILA}}$ consisted of original pretrained weights $\bm{\theta}$ and FILA-initialized low-rank adapter weights for each linear layer $\bm{\theta}_{\text{FILA}}=\{{\bm{A}}^{*}_{\ell},{\bm{B}}^{*}_{\ell}\}_{\ell=1}^{L}$, where $L$ represents the number of layers tuned via LoRA. Additionally, we incorporate GD, which utilizes the auxiliary retain set $\mathcal{D}_{r}$. The final loss function using both the proposed IHL and FILA is defined as follows:
@@
In practice, training (unlearning) for the LLM model is conducted by minimizing Equation 3 through stochastic gradient descent.
@@ 4 Experiments
Table 1: Evaluation results on reasoning and generative capabilities before and after unlearning samples from the TDEC dataset. For each model, we test three unlearning objectives (i.e. GA, GD, IHL) with full-parameter tuning, and four methods using LoRA with rank 16. IHL denotes replacing the GA loss term in GD with our IHL, and +FILA denotes initializing LoRA weights with FILA. Changes in performance after unlearning are in parentheses, with best results under LoRA-based unlearning in bold.
@@
GPT-Neo 125M Before - - 30.9 77.4 43.4 9.4 17.8
GA 100.0 17.2 1.0 27.4 39.9 (-3.5) 2.6 (-6.8) 577.8 (+560.0)
GD 4.6 0.7 24.9 42.4 (-1.0) 5.9 (-3.5) 54.2 (+36.4)
IHL 17.2 0.7 29.2 42.3 (-1.1) 10.3 (+0.9) 18.1 (+0.3)
GD 1.6 8.6 0.3 20.6 40.8 (-2.6) 2.5 (-6.9) 129.4 (+111.6)
IHL 11.4 0.4 22.7 41.9 (-1.5) 6.0 (-3.4) 32.9 (+15.1)
GD+FILA 7.4 1.2 27.4 42.0 (-1.4) 6.5 (-2.9) 89.5 (+71.7)
IHL+FILA 6.0 0.3 23.9 42.2 (-1.2) 10.1 (+0.7) 24.0 (+6.2)
GPT-Neo 1.3B Before - - 67.6 92.2 49.8 11.5 11.5
GA 100.0 13.8 1.9 30.4 49.7 (-0.1) 8.5 (-3.0) 15.8 (+4.3)
GD 12.8 2.2 30.9 48.4 (-1.4) 12.7 (+1.2) 10.8 (-0.7)
IHL 7.6 0.7 30.4 48.4 (-1.4) 12.5 (+1.0) 11.0 (-0.5)
GD 0.8 19.3 1.7 31.4 45.0 (-4.8) 9.7 (-1.8) 31.8 (+20.3)
IHL 20.0 1.7 44.6 47.1 (-2.7) 10.2 (-1.3) 14.9 (+3.4)
GD+FILA 7.8 1.9 23.2 44.2 (-5.6) 5.5 (-6.0) 54.5 (+43.0)
IHL+FILA 13.0 0.5 29.6 48.3 (-1.5) 12.1 (+0.6) 14.7 (+3.2)
GPT-Neo 2.7B Before - - 70.4 93.4 52.3 11.5 10.4
GA 100.0 10.8 1.6 31.0 51.9 (-0.4) 11.1 (-0.4) 17.9 (+7.5)
GD 8.0 0.7 28.3 51.8 (-0.5) 12.7 (+1.2) 17.9 (+7.5)
IHL 6.6 0.5 29.3 51.8 (-0.5) 12.9 (+1.4) 10.7 (+0.3)
GD 0.7 14.0 0.1 20.4 45.9 (-6.4) 6.7 (-4.8) 61.1 (+50.7)
IHL 17.8 0.0 26.7 49.6 (-2.7) 8.5 (-2.6) 22.2 (+11.8)
GD+FILA 6.8 1.6 28.9 44.8 (-7.5) 9.3 (-2.2) 68.7 (+58.3)
IHL+FILA 10.3 0.1 28.5 49.6 (-2.7) 10.7 (-0.8) 16.0 (+5.6)
@@ 4.1 Training Data Extraction Challenge
Experimental Setup.    The Training Data Extraction Challenge (TDEC) dataset (Carlini et al., 2021) consists of 20k examples from the Pile dataset (Gao et al., 2020) found to be easily extractable from a pretrained LLM. For each experiment, we randomly sample 32 sequences with 200 tokens to consist the forget set ${\mathcal{D}}_{f}$. For the retain set ${\mathcal{D}}_{r}$, we use the subset of WikiText (Merity et al., 2017) as it contains factual world knowledge that we wish to maintain after unlearning. We consider GPT-Neo 125M, 1.3B, and 2.7B pretrained on the Pile dataset as our base models, and unlearn ${\mathcal{D}}_{f}$ using five different forget sets. For this experiment, we use a fixed learning rate of 2e-4 and use LoRA adapters with rank $r=\{4,8,16,32\}$. For reasons we illustrate later in §4.2, we choose to apply LoRA on query and value layers in the attention module and two linear layers within feed-forward layers.
@@
Figure 3: Results from unlearning examples in the TDEC dataset on the GPT-Neo LLM family. Each row represents the performance averaged across datasets within each set of LLM capability tests: Reasoning (higher is better), Dialogue (higher is better), and Perplexity (lower is better). The circles and crosses represent successful and unsuccessful attempts, respectively, of unlearning a particular forget set ${\mathcal{D}}_{f}$. Solid lines indicate the performance of different methods averaged only across successful unlearning trials. The dashed lines indicate the base model performance prior to unlearning. GD leads to significant loss in performance and also fails to unlearn in some cases even with large LoRA ranks. Replacing the NCE loss in GD with IHL boosts retention of reasoning and generation capabilities, but still fails to unlearn in multiple cases. Running GD with FILA notably increases the rate of unlearning success, but at significant cost in overall performance. Using both IHL and FILA best minimizes post-unlearning performance degradation in all three aspects.
@@
Results.   Table 1 shows evaluation results using a fixed LoRA rank of 16 and Figure 3 shows analogous results using a different LoRA ranks. Our key findings are as follows. First, using GD not only meets the forgetting criteria in fewer epochs across all model sizes but it also preserves previously acquired knowledge (i.e., performance on Reasoning, Dialogue, and Pile) better than GA. Mainly for 125M and 1.3B models, GA causes a significant decline in generative performance, whereas GD partially mitigates this decline and improves both Dialogue F1 scores and the Pile perplexity. Second, we find that simply applying GD with LoRA fails to unlearn effectively across all model sizes. While enjoying great parameter-efficiency by tuning only about 0.7% to 1.6% of the total parameters when using LoRA, its application to unlearning with GD results in large loss in overall language capability, especially on generative tasks. Third, replacing GA in GD with IHL leads to performance gains in both full-parameter and LoRA-based unlearning, but requires larger number of epochs for successful unlearning than GD especially when confined to low-rank weight changes. However, this is resolved when IHL is used together with FILA, which significantly reduces the number of required epochs. Note that using FILA with GD also reduces the number of epochs required, but worsens downstream performance in many cases, as FILA essentially accelerates the divergent behavior of GA in GD. In essence, whether the ability of FILA to accelerate tuning also translates to benefits in downstream performance depends on the loss function being optimized. On the other hand, when paired with our bounded IHL, FILA leads to better retention of LLM performance, showcasing the strong synergy of our approach that enjoys not only the stability of IHL but also the speedup from FILA.
@@ 4.2 Analysis
Figure 4: Results from unlearning examples from TDEC dataset using LoRA with rank 32 to adapt sets of layers on GPT-Neo-125M. The marker shapes and colors are used similarly as in Figure 3. Based on the rate of unlearning success, tuning FFN layers (e.g., FFN, QVFFN) is more receptive to targeted knowledge removal compared to tuning attention layers (e.g., QV, QKVO).
@@ 4.3 Task of Fictitious Unlearning
Experimental Setup.  The Task of Fictitious Unlearning (TOFU) benchmark (Maini et al., 2024) is a synthetic dataset containing 20 question-answer pairs for each of 200 fictitious author profiles generated by GPT-4. The TOFU evaluation pipeline first finetunes a pretrained LLM on all QA pairs. Given this finetuned LLM that serves as our base model, our task is to unlearn all information regarding 1%, 5%, or 10% of the authors from the model. Note that we can obtain reference models finetuned only on the retain set (QA-pairs on 99%, 95%, or 90% of authors), with which we evaluate the Forget Quality of unlearned models by measuring the p-value from a Kolmogorov-Smirnov test. A high p-value indicates high distributional similarity between the unlearned model and the reference model, thus implying strong forgetting. To evaluate how well the model retains other information outside the forget set, we measure the Model Utility as the aggregation of the probability of correct answers, ROUGE-L scores, and Truth Ratio of correct answers vs. incorrect answers on questions from three datasets pertaining to the retain set of fictitious authors, real authors, and world facts.
Following the original paper of TOFU, we prepare two base models by finetuning Phi-1.5B and Llama2-7B on TOFU for 5 epochs with learning rates 2e-5 and 1e-5, respectively. We then unlearn with various methods using LoRA adapters of rank 4, 8, 16, or 32. While we mainly compare our methods against GD, we also compare IHL and FILA against existing unlearning methods such as KL (Maini et al., 2024), DPO (Rafailov et al., 2024), and NPO (Zhang et al., 2024), results from which can be found in Appendix D. For unlearning, we use a learning rate of 2e-4 if our base model is from Phi-1.5B and 1e-4 for Llama2-7B. All training procedures run 5 epochs with an effective batch size of 32 using the AdamW optimizer (Loshchilov & Hutter, 2019).
@@
Figure 5: TOFU results using Phi-1.5B and Llama2-7B models. Each row corresponds to unlearning a different forget set (1%, 5%, or 10%), and each column uses a distinct LoRA rank between 4 and 32. The relative size of markers represent the number of epochs. Ideally, the unlearning curves should start from the pretrained model ($\blacksquare$) and approach towards the reference model tuned on the retain set only ($\bigstar$) as unlearning progresses. Both GD and GD+FILA suffers from significant loss of model utility due to using GA for unlearning. Replacing GA with IHL largely retains model utility, then initializing LoRA adapters with FILA significantly boosts the unlearning efficiency of IHL.
@@
With regards to our proposed method, we find that replacing the NCE loss in GD with our IHL better retains model utility across all LoRA ranks and forget set sizes, as curves are more aligned straight-up towards the reference point with negligible shift in model utility. This stability comes at the cost of unlearning efficiency, however, as randomly initialized LoRA weights are unable to effectively represent weight changes required to decrease IHL. Nonetheless, initializing LoRA adapters with FILA largely alleviates this issue and significantly enhances unlearning efficiency of IHL by focusing gradient updates on parameters important to generating ${\mathcal{D}}_{f}$.